OpenStudy (anonymous):

If x+2=y and y^2-x^2=40 what is the value of x+1?

OpenStudy (anonymous):

if x + 2 = y that means x - y = 2 agree?

OpenStudy (anonymous):

oh yes! :)

OpenStudy (anonymous):

but then wat?

OpenStudy (anonymous):

@mathsuxx123

OpenStudy (loser66):

x=y-2 y^2-(y-2)^2=40 solve for y , I got y =11 therefor x =11-2 =9 --> x+1=10

OpenStudy (anonymous):

oh okay got it! Thank you so much :)

OpenStudy (anonymous):

@mathsuxx123 that is actually incorrect :) if x + 2 = y then x - y = -2

OpenStudy (anonymous):

@yummydum but @Loser66 ways works too

OpenStudy (anonymous):

@karolinarivas5 yes i know his way works best, i was just pointing out a mistake, irrelevant :)

OpenStudy (anonymous):

oh okay! thank you! :)

OpenStudy (anonymous):

yes @yummydum thank you for that correction